The perimeter of a rectangular Garden is 62 feet. The length is 1 foot more than twice the width . Find the dimension Of the garden
Answer by ikleyn(52787)

.
The perimeter of a rectangular Garden is 62 feet. The length is 1 foot more than twice the width . Find the dimension Of the garden
~~~~~~~~~~~~~~~~~

L = 2W + 1      ("The length is 1 foot more than twice the width.")

Perimeter = L + W + L + W = 62,   or

(2W + 1) + W + (2W+ 1) + W = 62.

6W + 2 = 62,

6W = 62 - 2,

6W = 60,

W = 10.   Then L = 2W + 1 = 2*10 + 1 = 21.